OpenStudy (anonymous):

is (1,5) a solution to the equation y=x-2

OpenStudy (anonymous):

To check if it is is true, we simply plug the values in and see if it works. So, in y=x-2, we put in x=1 and y=5 The equation becomes (5) = (1) - 2 Or, 5 = -1 Well this isn't correct, so the point is not a solution

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

solve the equation for x 3x = -4y a) x = -12 b) x = -4y/3 c) x = -12/y d) x = -3/4y

OpenStudy (whpalmer4):

@INeedSomeHelp! It is impolite to post your problem inside someone else's! Please do not do so.